OpenStudy (anonymous):

80 percent of the 160 4th grade girls had long hair, and 85 percent of the 120 5th grade girls had long hair. Which class had more girls with long hair, and by how many? A. 4th grade, 5 more B. 4th grade, 26 more C. 5th grade, 5 more D. 5th grade, 26 more

OpenStudy (lukecrayonz):

80% of 160 = 128 girls That's found by 160*0.8=128 85% of 120 = 102 girls 120*0.85=102

OpenStudy (anonymous):

@CashBoy27

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Just say B....

OpenStudy (anonymous):

80% of 160 = 128 girls U multiply 160 * 0.80 =128 85% of 120 = 102 girls 120 * 0.85=102